# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v2

@@ III.2 BdG equations in the generalized Gibbs ensemble
We can then identify $z_{\alpha}$ as the NG mode associated to the spontaneous symmetry breaking in a particular application of the general result of Eq. (7). In turn, each NG mode $z_{\alpha}$ is paired with a Gibbs mode $z_{A}$, obtained by deriving the GPG equation with respect to the corresponding conserved charge $Q_{\alpha}$ [see Eq. (26)],
@@
In the following, for simplicity, we assume that there are no dynamical instabilities (i.e., there are no complex frequency modes) and that the only modes with zero frequency are the NG modes. It is immediate to show that one can separate the BdGG modes into two orthogonal sectors: the Goldstone-Gibbs sector formed by the paired modes $\{z_{a}\}=\{z_{\alpha},z_{A}\}$, and the regular Bogoliubov sector formed by the eigenmodes $\{z_{n}\}$,
@@
$\partial_{b}Q_{\alpha}=\int\mathrm{d}\mathbf{x}~{}\partial_{b}\Psi^{*}T_{\alpha}\Psi+\partial_{b}\Psi(T_{\alpha}\Psi)^{*}=-i(z_{\alpha}|z_{b}).$ (48)
@@
is nothing else than the usual Berry connection evaluated for the GPG wavefunction, which here depends on the set of conserved charges $\mathbf{Q}$ (we recall they are not intrinsic parameters of the original Hamiltonian). We then immediately identify $F=dW$, with $W_{A}=i\braket{\Psi_{0}}{\partial_{A}\Psi_{0}}$  [^2] Strictly speaking, it should be $W_{A}=i(\braket{\Psi_{0}}{\partial_{A}\Psi_{0}}-\braket{\partial_{A}\Psi_{0}}{\Psi_{0}})/2$ to ensure that $W_{A}$ is real since $\partial_{A}\braket{\Psi_{0}}{\Psi_{0}}=\partial_{A}N\neq 0$.. As a result of this analogy, we denote $F_{AB}$ as the Berry-Gibbs connection. Its main differences with the standard Berry connection are i) $\braket{\Psi_{0}}{\Psi_{0}}$ does depend on the connection parameters (specifically, $\braket{\Psi_{0}}{\Psi_{0}}=N$); and ii) we can perform generalized gauge transformations, which go beyond the usual phase transformations $\Psi_{0}\to e^{-i\theta(\mathbf{Q})}\Psi_{0}$, under which $F_{AB}$ is not invariant.
The Berry-Gibbs connection can be extended to the complete Goldstone-Gibbs manifold with coordinates $x^{a}=(x^{\alpha},x^{A})=(\boldsymbol{\alpha},\mathbf{Q})$ as
@@
$I_{n}$ being the $n\times n$ identity matrix. The off-diagonal blocks are determined by Eq. (49). In this work, we have assumed that the generators of the broken symmetries commute between themselves. This implies that $F_{\alpha\beta}=0$, Eq. (47), and that one can always set $F_{AB}=0$ by means of a generalized gauge transformation, as discussed above. Actually, within the complete Goldstone-Gibbs manifold, a generalized gauge transformation is nothing else than a change of coordinates
@@ III.3 Quantization of the Nambu-Goldstone modes
The symmetry transformations on the field operator are now implemented as
@@
Now, we expand the field operator around its GPG expectation value, $\hat{\Psi}^{\prime}(\mathbf{x},t)=\Psi_{0}(\mathbf{x})+\hat{\varphi}(\mathbf{x},t)$, recovering at linear order the quantum time-dependent BdGG equations (42),
@@
$\hat{X}^{\alpha}$$\equiv$ $\hat{\gamma}^{\alpha}=i(z_{A}|\hat{\Phi})$ (73)
$\hat{P}^{A}$$\equiv$ $\hat{\gamma}^{A}=-i(z_{\alpha}|\hat{\Phi})$
@@
$\hat{X}_{A}(t)=\hat{X}_{A}+(\partial_{A}\lambda_{B})\hat{P}^{B}t=\hat{X}_{A}+(\partial_{B}\lambda_{A})\hat{P}^{B}t$ (75)
The physics behind this equation is very simple. After quantizing, each conserved charge is expanded up to linear order in the field fluctuations in the same fashion as Eq. (48),
@@
The quantum fluctuations of the velocities are then translated into a ballistic motion for the quantum amplitude of the NG modes, Eq. (75). Another way to put it is that the modes $z_{A}$ are not eigenmodes of the BdGG equations. Instead, from Eq. (24), we have that
@@
The second term is the usual Bogoliubov contribution while the first one is the Goldstone-Gibbs contribution. The latter is quadratic in the fluctuations of the conserved charges, then behaving as effective momenta, and does not depend on the NG amplitudes, since the Hamiltonian is invariant under the symmetry transformations. In particular, if the solution is energetically stable, then $\partial^{2}_{AB}E$ is a positive definite form. In fact, one can always diagonalize $\partial^{2}_{AB}E$ by performing an appropriate rotation in the Goldstone-Gibbs sector, $\partial^{2}_{AB}E=M_{A}^{-1}\delta_{AB}$, where $M_{A}$ behaves as the effective mass of each mode. This rotation yields a new set of independent Goldstone-Gibbs modes $\{z^{\prime}_{\alpha},z^{\prime}_{A}\}$, whose amplitudes $\{\hat{X}^{\prime\alpha},\hat{P}^{\prime A}\}$ satisfy
@@
Thus, the matrix $\partial^{2}_{AB}E$, computed from thermodynamical considerations, determines how the different NG modes hybridize in a dynamical context. Moreover, as NG modes are the zero-frequency limit of some branches from the regular Bogoliubov sector, the structure of the hybrid Goldstone-Gibbs modes $\{z^{\prime}_{\alpha},z^{\prime}_{A}\}$ also describes their low-frequency regime.
The quantization of the Goldstone-Gibbs modes in the usual GP context is simply retrieved by first setting all Lagrange multipliers to zero except for the chemical potential $\lambda^{\theta}=\mu$, so the only independent conserved charge is the particle number $N$. With the resulting GP solution, one computes the remaining charges $Q_{\alpha}$ to assemble the vector $\mathbf{Q}=\mathbf{Q}_{0}$. Once known, the full GPG equation is solved as a function of $\mathbf{Q}$, from where the Berry-Gibbs connection $F_{AB}$ and the matrix $\partial^{2}_{AB}E$ can be evaluated at the point $\mathbf{Q}_{0}$ characterizing the GP solution.
@@ IV Cnoidal wave in a superfluid
Technical details behind the calculations of this section as well as basic properties of elliptic functions are presented in Appendix A.
@@
We have explicitly evaluated the extended Berry-Gibbs connection $F_{ab}=i(z_{a}|z_{b})$, recovering the predicted symplectic form, $F_{ab}=-\Omega_{ab}$. Specifically, in our gauge choice, the Berry-Gibbs connection automatically vanishes without the need of further generalized gauge transformations, $F_{NP}=F_{PN}=0$ [see Eq. (188) and ensuing discussion]. In this example, a generalized gauge transformation amounts to add global phase and spatial shifts that depend on the particle number and momentum, $\theta=\theta(N,P),~{}x_{0}=x_{0}(N,P)$. The quantum amplitudes of the NG modes can be then understood as the quantum fluctuations of the global phase and position of the cnoidal wave, $\hat{X}^{\theta}=\delta\hat{\theta}$, $\hat{X}^{x}=\delta\hat{x}_{0}$, while their momenta describe total particle number and momentum fluctuations, $\hat{P}^{N}=\delta\hat{N}$, $\hat{P}^{P}=\delta\hat{P}$.
Regarding thermodynamics, in all numerical cases we have found that the matrix $\partial^{2}_{AB}E$ is both non-diagonal and not positive definite. This reflects, respectively, that both phase and crystalline modes hybridize in the low-frequency limit, and the presence of a negative branch in the regular Bogoliubov sector as the true energy minimum is always an homogeneous GP solution Martone et al. (2021). In fact, the results of Ref. Martone et al. (2021), derived within the standard mean-field GP approach, correspond to setting $v=0$ instead of fixing the total momentum $P$, and taking the limit $L\to\infty$. However, even for $v=0$, the full GPG equation is still needed for the computation of $\partial^{2}_{AB}E$ and the subsequent characterization of the NG modes [see discussion after Eq. (81)].
@@
This analysis suggests that certain time-periodic systems that do not genuinely break time-translational symmetry can be misidentified as continuous time crystals, with the time periodicity arising from constant motion along a closed orbit spawned by other broken generators.
@@ V Variational Floquet-Goldstone Theorem: Floquet-Nambu-Goldstone modes
In order to study spontaneous symmetry breaking of time-translation symmetry, we first analyze Floquet systems described by periodic Hamiltonians, $\hat{H}(t+T)=\hat{H}(t)$, explicitly breaking this symmetry. The periodic driving may not only be applied in the external potential Lignier et al. (2007); Kierig et al. (2008), but also in the interacting Gaul et al. (2011); Rapp et al. (2012) and even in the kinetic term Stöferle et al. (2004). In a many-body system whose dynamics can be described by a variational ansatz of the form (4), a periodic Hamiltonian is translated into
@@
with $M^{i}_{j}(t)$ also periodic. As a linear periodic equation, it admits Floquet solutions of the form $\delta\bar{X}^{i}(t)=u^{i}_{\varepsilon}(t)e^{-i\varepsilon t}$, with $u^{i}_{\varepsilon}(t+T)=u^{i}_{\varepsilon}(t)$ and $\varepsilon$ the quasi-energy. The corresponding eigenvalue equation reads
@@
It was recently shown de Nova and Sols (2022) that, due to many-body interactions, spontaneous Floquet states can emerge even in the absence of external periodic driving as self-consistent Floquet solutions (109) to the original time-independent problem (4),
@@
It is important to remark that here the periodicity of $H^{i}_{j}(\bar{X}(t))$ stems solely from that of $\bar{X}(t)$, and is spontaneously determined by the solution itself and not externally imposed, in contrast to Eq. (110). The linear problem is also described by Floquet modes, as in Eq. (112), but once more the periodicity of the matrix $M^{i}_{j}(t)$ stems just from that of $\bar{X}(t)$. Similarly, spontaneous symmetry breaking is translated into the emergence of a FNG mode. However, a spontaneous Floquet state also breaks time-translation symmetry as $\bar{X}(t+t_{0})$ is as well a solution of Eq. (114). This implies that $\partial_{t}\bar{X}(t)$ is a FNG mode,
@@ VI Simultaneous symmetry breaking in spontaneous Floquet states
When inserted in the usual time-dependent GP equation, taking $\Psi(\mathbf{x},0)=\Psi_{0}(\mathbf{x},0)$ yields
@@
The structure of these equations suggests to gather hereafter the symmetries $\mathbf{T}$ (which we recall do not involve time by assumption) with the time-translation symmetry as
@@
with $M_{0}(t)$ the periodic operator resulting from the substitution $\Psi_{0}(\mathbf{x})\to\Psi_{0}(\mathbf{x},t)$ in Eq. (38). Thus, the modes $z_{\alpha}(t)$ are zero quasi-energy modes, which we can identify as the FNG modes associated to the spontaneous symmetry breaking by virtue of Eq. (113). The Floquet spectrum is completed by the regular Bogoliubov sector
$[M-i\partial_{t}]z_{\nu,\varepsilon}(\mathbf{x},t)=\varepsilon z_{\nu,\varepsilon}(\mathbf{x},t),~{}z_{\nu,\varepsilon}(\mathbf{x},t+T)=z_{\nu,\varepsilon}(\mathbf{x},t),$ (121)
with $\varepsilon$ the quasi-energy (defined modulo $\omega=2\pi/T$) and $\nu$ a discrete index labeling the solution. With the help from the conventional Floquet theory and from the results of Sec. III.2, it is shown that the orthogonality relations (57) still hold, namely
$(z_{\nu,\varepsilon}|z_{\nu^{\prime},\varepsilon^{\prime}})=\delta_{\nu\nu^{\prime}}\delta_{\varepsilon\varepsilon^{\prime}},~{}(z_{a}|z_{\nu,\varepsilon})=0,~{}(z_{a}|z_{b})=i\Omega_{ab}.$ (122)
@@
$+$ $\sum_{\nu,\varepsilon}\hat{\gamma}_{\nu,\varepsilon}(t)z_{\nu,\varepsilon}(\mathbf{x},t)+\hat{\gamma}^{\dagger}_{\nu,\varepsilon}(t)\bar{z}_{\nu,\varepsilon}(\mathbf{x},t).$
@@
This results in the same time dependence for the quantum amplitudes as in the stationary case, Eq. (III.3), namely:
$\hat{\gamma}_{\nu,\varepsilon}(t)=\hat{\gamma}_{\nu,\varepsilon}e^{-i\varepsilon t},\hat{P}^{A}(t)=\hat{P}^{A},\hat{X}^{\alpha}(t)=\hat{X}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}^{B}t.$ (125)

### v3

@@ III.2 Nambu-Goldstone modes in the generalized Gibbs ensemble
We can then identify $z_{\alpha}$ as the NG mode associated to the spontaneous symmetry breaking in a particular application of the general result of Eq. (7). In turn, each NG mode $z_{\alpha}$ is paired with a Gibbs mode $z_{A}$, obtained by taking derivative of the GPG equation with respect to the corresponding conserved charge $Q_{\alpha}$ [see Eq. (26)],
@@
In the following, for simplicity, we assume that there are no dynamical instabilities (i.e., there are no complex frequency modes) and that the only modes with zero frequency are the NG modes. It is immediate to show that one can separate the BdGG modes into two orthogonal sectors: the Goldstone-Gibbs sector formed by the paired modes $\{z_{a}\}=\{z_{\alpha},z_{A}\}$, and the regular Bogoliubov sector formed by the finite-frequency eigenmodes $\{z_{n}\}$,
@@ III.3 Berry-Gibbs connection
$\partial_{b}Q_{\alpha}=\int\mathrm{d}\mathbf{x}~{}(T_{\alpha}\Psi_{0})^{*}\partial_{b}\Psi_{0}+(\partial_{b}\Psi_{0})^{*}T_{\alpha}\Psi_{0}=-i(z_{\alpha}|z_{b}).$ (48)
@@
is nothing else than the usual Berry curvature evaluated for the GPG wavefunction, which here depends on the set of conserved charges $\mathbf{Q}$ (we recall that they are not intrinsic parameters of the original Hamiltonian). We then immediately identify $F=dW$, with $W_{A}=i\braket{\Psi_{0}}{\partial_{A}\Psi_{0}}$  [^2] Strictly speaking, it should be $W_{A}=i(\braket{\Psi_{0}}{\partial_{A}\Psi_{0}}-\braket{\partial_{A}\Psi_{0}}{\Psi_{0}})/2$ to ensure that $W_{A}$ is real since $\partial_{A}\braket{\Psi_{0}}{\Psi_{0}}=\partial_{A}N\neq 0$.. As a result of this analogy, we denote the vector $W$ as the Berry-Gibbs connection and $F_{AB}$ as the Berry-Gibbs curvature. Its main differences with the standard Berry connection are i) $\braket{\Psi_{0}}{\Psi_{0}}$ does depend on the connection parameters (specifically, $\braket{\Psi_{0}}{\Psi_{0}}=N$); and ii) we can perform generalized gauge transformations, which go beyond the usual phase transformations $\Psi_{0}\to e^{-i\theta(\mathbf{Q})}\Psi_{0}$, under which $F_{AB}$ is not invariant.
The Berry-Gibbs curvature can be extended to the complete Goldstone-Gibbs manifold with coordinates $x^{a}=(x^{\alpha},x^{A})=(\boldsymbol{\alpha},\mathbf{Q})$ as
@@
$I_{n}$ being the $n\times n$ identity matrix. The off-diagonal blocks are determined by Eq. (49). In this work, we have assumed that the generators of the broken symmetries commute between themselves. This implies that $F_{\alpha\beta}=0$, Eq. (47), and that one can always set $F_{AB}=0$ by means of a generalized gauge transformation, as discussed above. Actually, within the complete Goldstone-Gibbs manifold, a generalized gauge transformation is just a change of coordinates
@@ III.4 Quantization
In the case of a fermionic field operator, the commutator is replaced by an anticommutator. The symmetry transformations on the field operator can be implemented as
@@
These results are general and exact, easily translatable to an arbitrary many-body system. In the specific case of a condensate, we can expand the field operator around the GPG expectation value, $\hat{\Psi}^{\prime}(\mathbf{x},t)=\Psi_{0}(\mathbf{x})+\hat{\varphi}(\mathbf{x},t)$, recovering at linear order the quantum time-dependent BdGG equations (42),
@@
$\hat{X}^{\alpha}$$\equiv$ $\hat{\gamma}^{\alpha}=i(z_{A}|\hat{\Phi}),$ (73)
$\hat{P}^{A}$$\equiv$ $\hat{\gamma}^{A}=-i(z_{\alpha}|\hat{\Phi}),$
@@
$\hat{X}_{A}(t)=\hat{X}_{A}+(\partial_{A}\lambda_{B})\hat{P}^{B}t=\hat{X}_{A}+(\partial_{B}\lambda_{A})\hat{P}^{B}t.$ (75)
The physics behind this equation is very simple. After quantization, each conserved charge is expanded up to linear order in the field fluctuations in the same fashion as Eq. (48),
@@
The quantum fluctuations of the velocities are then translated into a ballistic motion for the quantum amplitude of the NG modes, arriving at Eq. (75). Another way to put it is that the modes $z_{A}$ are not eigenmodes of the BdGG equations. Instead, from Eq. (24), we have that
@@
The second term is the usual Bogoliubov contribution while the first one is the Goldstone-Gibbs contribution. The latter is quadratic in the fluctuations of the conserved charges, then behaving as effective momenta, and does not depend on the NG amplitudes, since the Hamiltonian is invariant under the symmetry transformations. In particular, if the solution is energetically stable, then $\partial^{2}_{AB}E$ is a positive definite form. In fact, one can always diagonalize $\partial^{2}_{AB}E$ by performing an appropriate rotation in the Goldstone-Gibbs sector, $\partial^{2}_{AB}E=M_{A}^{-1}\delta_{AB}$, where $M_{A}$ plays the role of an effective mass. This rotation yields a new set of independent Goldstone-Gibbs modes $\{z^{\prime}_{\alpha},z^{\prime}_{A}\}$, whose amplitudes $\{\hat{X}^{\prime\alpha},\hat{P}^{\prime A}\}$ satisfy
@@
Thus, the matrix $\partial^{2}_{AB}E$, computed from thermodynamical considerations, determines how the different NG modes hybridize in a dynamical context. Moreover, as the NG modes are the zero-frequency limit of some branches from the regular Bogoliubov sector, the structure of the hybrid Goldstone-Gibbs modes $\{z^{\prime}_{\alpha},z^{\prime}_{A}\}$ also describes their low-frequency regime.
The quantization of the Goldstone-Gibbs modes in the usual GP context is simply retrieved by first setting all Lagrange multipliers to zero except for the chemical potential $\lambda^{\theta}=\mu$, so the only independent conserved charge is the particle number $N$. With the resulting GP solution, one computes the remaining charges $Q_{\alpha}$ to assemble the vector $\mathbf{Q}=\mathbf{Q}_{0}$. Once known, the full GPG equation is solved as a function of $\mathbf{Q}$, from where the Berry-Gibbs curvature $F_{AB}$ and the matrix $\partial^{2}_{AB}E$ can be evaluated at the point $\mathbf{Q}_{0}$ characterizing the GP solution.
@@ IV Cnoidal wave in a superfluid
Technical details behind the calculations of this section as well as basic properties of elliptic functions are presented in the Appendix.
@@
We have explicitly evaluated the extended Berry-Gibbs curvature $F_{ab}=i(z_{a}|z_{b})$, recovering the predicted symplectic form, $F_{ab}=-\Omega_{ab}$. Specifically, in our gauge choice, the Berry-Gibbs curvature automatically vanishes without the need of further generalized gauge transformations, $F_{NP}=F_{PN}=0$ [see Eq. (191) and ensuing discussion]. In this example, a generalized gauge transformation amounts to add global phase and spatial shifts that depend on the particle number and momentum, $\theta=\theta(N,P),~{}x_{0}=x_{0}(N,P)$. The quantum amplitudes of the NG modes can be then understood as the quantum fluctuations of the global phase and position of the cnoidal wave, $\hat{X}^{\theta}=\delta\hat{\theta}$, $\hat{X}^{x}=\delta\hat{x}_{0}$, while their momenta describe total particle number and momentum fluctuations, $\hat{P}^{N}=\delta\hat{N}$, $\hat{P}^{P}=\delta\hat{P}$.
Regarding thermodynamics, in all numerical cases we have found that the matrix $\partial^{2}_{AB}E$ is both non-diagonal and not positive definite. This reflects, respectively, that both phase and crystalline modes hybridize in the low-frequency limit, and the presence of a negative-energy branch in the regular Bogoliubov sector as the true energy minimum is always an homogeneous GP solution Martone et al. (2021). In fact, the results of Ref. Martone et al. (2021), derived within the standard mean-field GP approach, correspond to setting $v=0$ instead of fixing the total momentum $P$, and taking the limit $L\to\infty$. However, even for $v=0$, the full GPG equation is still needed for the computation of $\partial^{2}_{AB}E$ and the subsequent characterization of the NG modes [see discussion after Eq. (81)].
@@
This analysis suggests that certain time-periodic systems that do not genuinely break time-translational symmetry can be misidentified as continuous time crystals, with the time periodicity simply arising from constant motion along a closed orbit spawned by other broken generators.
@@ V Variational Floquet-Goldstone Theorem: Floquet-Nambu-Goldstone modes
In order to study spontaneous symmetry breaking of time-translation symmetry, we first analyze Floquet systems, which are described by periodic Hamiltonians explicitly breaking this symmetry, $\hat{H}(t+T)=\hat{H}(t)$. The periodic driving may not only be applied in the external potential Lignier et al. (2007); Kierig et al. (2008), but also in the interacting Gaul et al. (2011); Rapp et al. (2012) and even in the kinetic term Stöferle et al. (2004). In a many-body system whose dynamics can be described by a variational ansatz of the form (4), a periodic Hamiltonian is translated into
@@
with $M^{i}_{j}(t+T)=M^{i}_{j}(t)$. As a linear periodic equation, it admits Floquet solutions of the form $\delta\bar{X}^{i}(t)=u^{i}_{\varepsilon}(t)e^{-i\varepsilon t}$, with $u^{i}_{\varepsilon}(t+T)=u^{i}_{\varepsilon}(t)$ and $\varepsilon$ the quasi-energy. The corresponding eigenvalue equation reads
@@
It was recently shown de Nova and Sols (2022) that, due to many-body interactions, Floquet states can spontaneously emerge even in the absence of external periodic driving as self-consistent Floquet solutions (109) to the original time-independent problem (4),
@@
It is important to remark that here the periodicity of $H^{i}_{j}(\bar{X}(t))$ stems solely from that of $\bar{X}(t)$, and is spontaneously determined by the solution itself and not externally imposed, in contrast to Eq. (110). The linear problem is also described by Floquet modes, as in Eq. (112), but once more the periodicity of the matrix $M^{i}_{j}(t)$ stems just from that of $\bar{X}(t)$. Similarly, spontaneous symmetry breaking is translated into the emergence of FNG modes. However, a spontaneous Floquet state also breaks time-translation symmetry as $\bar{X}(t+t_{0})$ is as well a solution of Eq. (114). This implies that $\partial_{t}\bar{X}(t)$ is a FNG mode,
@@ VI Simultaneous symmetry breaking in spontaneous Floquet states
When inserted in the usual time-dependent GP equation (22), taking $\Psi(\mathbf{x},0)=\Psi_{0}(\mathbf{x},0)$ yields
@@
Due to time-translation symmetry, if $\Psi_{0}(\mathbf{x},t)$ is a solution of this equation, then $\Psi_{0}(\mathbf{x},t+t_{0})$ is also a solution. The structure of these equations suggests to gather hereafter the symmetries $\mathbf{T}$ (which we recall do not involve time by assumption) with the time-translation symmetry as
@@
where $M_{0}(t)$ is the periodic operator resulting from the substitution $\Psi_{0}(\mathbf{x})\to\Psi_{0}(\mathbf{x},t)$ in Eq. (38). Thus, the modes $z_{\alpha}(t)$ are zero quasi-energy modes, which we can identify as the FNG modes associated to the spontaneous symmetry breaking by virtue of Eq. (113). The Floquet spectrum is completed by the regular Bogoliubov sector
$[M_{0}(t)-i\partial_{t}]z_{\varepsilon,\nu}=\varepsilon z_{\varepsilon,\nu},$ (121)
with $z_{\varepsilon,\nu}(\mathbf{x},t+T)=z_{\varepsilon,\nu}(\mathbf{x},t)$, $\varepsilon$ the quasi-energy (defined modulo $\omega=2\pi/T$), and $\nu$ a discrete index labeling the solution. With the help from the conventional Floquet theory and from the results of Sec. III.2, it is shown that the orthogonality relations (57) still hold:
$(z_{\varepsilon,\nu}|z_{\varepsilon^{\prime},\nu^{\prime}})=\delta_{\varepsilon\varepsilon^{\prime}}\delta_{\nu\nu^{\prime}},~{}(z_{a}|z_{\varepsilon,\nu})=0,~{}(z_{a}|z_{b})=i\Omega_{ab}.$ (122)
@@
$+$ $\sum_{\varepsilon,\nu}\hat{\gamma}_{\varepsilon,\nu}(t)z_{\varepsilon,\nu}(\mathbf{x},t)+\hat{\gamma}^{\dagger}_{\varepsilon,\nu}(t)\bar{z}_{\varepsilon,\nu}(\mathbf{x},t).$
@@
This results in the same time dependence for the quantum amplitudes as in the stationary case, Eq. (III.4), namely:
$\hat{\gamma}_{\varepsilon,\nu}(t)=\hat{\gamma}_{\varepsilon,\nu}e^{-i\varepsilon t},\hat{P}^{A}(t)=\hat{P}^{A},\hat{X}^{\alpha}(t)=\hat{X}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}^{B}t.$ (125)
@@
Figure 1: Schematic representation of the $(t,\phi)$ formalism, where $t$ parametrizes the longitudinal axis of the cylinder and $\phi$ its polar angle. The dynamics for the field $\Psi(\mathbf{x},\phi,t)$ is solved over the whole cylinder surface while the physical field $\Psi(\mathbf{x},t)=\Psi(\mathbf{x},\phi_{0}+\omega t,t)$ is restricted to the spiral trajectory given by the solid black line. The red circles correspond to the constant sections $t=0$ and $t=T$.